# A four-operator splitting algorithm for nonconvex and nonsmooth optimization

### v1

@@ 4.1 Nonnegative and low-rank matrix completion
The first experiment we consider is nonnegative matrix completion/factorization (Lee and Seung, 1999), whose goal is to recover missing entries of a partially observed matrix using nonnegative entries, and we impose the additional structure requirement that the recovered matrix should be low-rank. Given a data matrix $M\in{\rm I}\!{\rm R}^{m\times n}$, we let $\Omega$ denote the entries $(i,j)$ that are observed and $P_{\Omega}$ denote the associated projection operation such that
$P_{\Omega}(X)_{i,j}=\begin{cases}X_{i,j}&\text{ if }(i,j)\in\Omega,\\ 0&\text{ if }(i,j)\notin\Omega.\end{cases}$
For this task, we consider Eq. 1.1 with
$f(X)=\frac{\lambda_{1}}{2}\min_{Y\in{\rm I}\!{\rm R}^{m\times n}_{+}}\|X-Y\|^{2}_{F},\quad g(X)=\lambda_{2}\|X\|_{*},\quad h(X)=\frac{1}{2}\|P_{\Omega}(X-M)\|^{2}_{F},\quad p(X)\equiv 0,$ (4.1)
@@ 4.2 Cardinality-constrained problems
Our next experiment considers a least-square problem with a penalized form of the cardinality constraint $\|x\|_{0}\leq k$ for some given positive integer $k$. As argued in (Gotoh et al., 2018), this constraint is equivalent to
$\|x\|_{1}-\|x\|_{(k)}=0,$ (4.2)
where $\|x\|_{(k)}$ is the Ky Fan $k$-norm that computes the sum of the top $k$ largest elements of the element-wise absolute value of $x$, and we follow Gotoh et al. (2018) to take $\|x\|_{1}-\|x\|_{(k)}$ as a penalty in the objective function instead of enforcing the hard constraint. We also follow a common convention in machine learning to add in a small Tikhonov regularization of the form $\|x\|_{2}^{2}$ to improve the problem condition. This leads to the following setting for Eq. 1.1.
@@
Given the existence of the $p$ term in this experiment, PG is not applicable, but by again treating $(f+h)$ as the smooth term, we can use the proximal DC algorithm described in Eq. 3.6. Our experiment uses publicly available real-world data listed in Table 2.333Downloaded from https://www.csie.ntu.edu.tw/~cjlin/libsvmtools/datasets/. In this experiment, we set $\lambda_{1}=0.01$, $\lambda_{2}=0.005$, and $k=\lfloor n/10\rfloor$. As the requirement Eq. 3.48 does not hold for any $\tau\geq 2$, for our method we use the same range of $\tau$ as the previous experiment. All algorithms are run with a cap of $100,000$ iterations. The results are shown in Table 3. We can see that with appropriately selected $\tau$, our method performs similarly to proximal DC on ijcnn1, and is much more efficient than it on all the other datasets.
@@
Table 3: Comparison between proximal DC and our method for cardinality-constrained least-square problems. * denotes that the maximum iteration of $100000$ is reached. $\tau=t$ denotes our method with the specified $\tau$.

### v2

@@ 4.1 Nonnegative and low-rank matrix completion
The first experiment we consider is nonnegative matrix completion/factorization (Lee and Seung, 1999), whose goal is to recover missing entries of a partially observed matrix using nonnegative entries, and we impose the additional structure requirement that the recovered matrix should be low-rank. Given a data matrix $M\in{\rm I}\!{\rm R}^{m\times n}$, we let $\Omega$ denote the entries $(i,j)$ oberserved and $P_{\Omega}$ denote the associated projection operation such that $P_{\Omega}(X)_{i,j}$ outputs $X_{i,j}$ if $(i,j)\in\Omega$ and 00 otherwise. We consider Eq. 1.1 with $p(X)\equiv 0$ and
$f(X)=\frac{\lambda_{1}}{2}\min_{Y\in{\rm I}\!{\rm R}^{m\times n}_{+}}\|X-Y\|^{2}_{F},\quad g(X)=\lambda_{2}\|X\|_{*},\quad h(X)=\frac{1}{2}\|P_{\Omega}(X-M)\|^{2}_{F},$
@@ 4.2 Cardinality-constrained problems
Our next experiment considers a least-square problem with a penalized form of the cardinality constraint $\|x\|_{0}\leq k$ for some given positive integer $k$. As argued in (Gotoh et al., 2018), this constraint is equivalent to $\|x\|_{1}-\|x\|_{(k)}=0$, where $\|x\|_{(k)}$ is the Ky Fan $k$-norm that computes the sum of the top $k$ largest elements of the element-wise absolute value of $x$, and we follow Gotoh et al. (2018) to take $\|x\|_{1}-\|x\|_{(k)}$ as a penalty in the objective function instead of enforcing the hard constraint. We also follow a common convention in machine learning to add in a small Tikhonov regularization of the form $\|x\|_{2}^{2}$ to improve the problem condition. This leads to the following setting for Eq. 1.1.
@@
Given the existence of the $p$ term in this experiment, PG is not applicable, but by again treating $(f+h)$ as the smooth term, we can use the proximal DC algorithm described in Eq. 3.6. Our experiment uses publicly available real-world data [^3] Downloaded from https://www.csie.ntu.edu.tw/~cjlin/libsvmtools/datasets/. listed in Table 2. In this experiment, we set $\lambda_{1}=0.01$, $\lambda_{2}=0.005$, and $k=\lfloor n/10\rfloor$. As Eq. 3.49 does not hold for any $\tau\geq 2$, for our method we use the same range of $\tau$ as the previous experiment. All algorithms are run with a cap of $100,000$ iterations. The results are shown in Table 3. We can see that with appropriately selected $\tau$, our method performs similarly to proximal DC on ijcnn1, and is much more efficient than it on all the other datasets.
@@
Table 3: Comparison between proximal DC (denoted by PDC) and our method for Eq. 4.1 * denotes that the maximum iteration of $100000$ is reached. $\tau=t$ denotes our method with the specified $\tau$.